# Generative LiDAR Editing with Controllable Novel Object Layouts

### v1

@@ Abstract
We propose a framework to edit real-world Lidar scans with novel object layouts while preserving a realistic background environment. Compared to the synthetic data generation frameworks where Lidar point clouds are generated from scratch, our framework focuses on new scenario generation in a given background environment, and our method also provides labels for the generated data. This approach ensures the generated data remains relevant to the specific environment, aiding both the development and the evaluation of algorithms in real-world scenarios. Compared with novel view synthesis, our framework allows the creation of counterfactual scenarios with significant changes in the object layout and does not rely on multi-frame optimization. In our framework, the object removal and insertion are supported by generative background inpainting and object point cloud completion, and the entire pipeline is built upon spherical voxelization, which realizes the correct Lidar projective geometry by construction. Experiments show that our framework generates realistic Lidar scans with object layout changes and benefits the development of Lidar-based self-driving systems.
@@ I Introduction
Data is a key factor in the realization of self-driving. Collecting a huge amount of data with good diversity is necessary for understanding the corner cases, which are low-frequency and high-risk real-world situations. Both the development and the testing of self-driving systems could benefit from data with good long-tail coverages. However, the low-frequency nature means that it is inherently expensive to collect corner-case data, and as progress is being made on relatively easier and more frequent cases, capturing even rarer events only becomes an even more challenging task. That’s why the research community has a growing interest in data generation, which allows for acquiring new data at a lower cost and potentially higher diversity compared with real-world data collection.
Early efforts of data synthesis rely on gaming engines and graphical rendering. This approach generates data from highly configurable virtual environments, but the scalability is limited, and the domain gap to the real data is non-negligible. Recent progress in generative modeling makes end-to-end data generation possible, largely improving the scalability and realism. However, users generally have less control over the specifics of the generated data. Another notable trend is novel-view synthesis from real-world sequential observations, allowing the resimulation of a real-world scenario from different viewpoints, which helps analyze and improve the algorithms. However, the variation flexibility of the re-simulated data is limited by the visible signal.
In this paper, we propose to solve the task of scene editing with controllable novel object layouts, with a specific focus on Lidar scans—a widely used and reliable sensory modality in autonomous vehicles. This new task enables us to freely remove and insert objects with arbitrary poses in a real Lidar scan, not limited by the visible information. On the other hand, the generated data is associated with a real environment, building a more direct connection to real-world performance. An illustration of this task is shown in Fig. 1. Leveraging recent progress in conditional lidar data generation and object point cloud completion, we build a lightweight framework without scene reconstruction, neural rendering, or end-to-end generation from scratch. We bake the prior knowledge of Lidar projective geometry into the framework through spherical voxelization so that the occlusion relationship and the scanning pattern are correct by construction. Experiments show that our framework enables flexible, controllable, and realistic Lidar scene editing, which is beneficial for both the development and testing of self-driving algorithms.
@@ II Related Work
We will review the literature related to Lidar data generation. Most work did not directly address the same problem as targeted in this paper, but a lot of them are relevant and prepared the techniques that are useful in our problem.
@@ II-A Lidar simulation
Lidar simulation refers to the approach of building a 3D environment with 3D object assets and casting rays on the 3D environment to create point clouds through ray-surface intersection [1]. While graphics-based simulation engines [2, 3, 4] facilitate the rendering, the requirement of pre-built 3D assets limits the scalability, and the rendering and the quality of the 3D assets both contribute to the domain gap compared with real data. [5] constructed the 3D assets from real-world scans to avoid manual asset creation and used learning-based postprocessing to reduce the domain gap. Later work [6] further investigated various Lidar phenomena that could improve the fidelity of Lidar simulators.
@@ II-B Lidar novel-view synthesis
The novel-view synthesis (NVS) of Lidar point clouds and Lidar simulation share the same conceptual procedure of 3D construction and rendering, but today’s Lidar NVS realizes these two steps in a unified learnable framework called neural rendering. The learned implicit geometry-appearance model and the differentiable rendering are both optimized towards reproducing more realistic synthesis, allowing a smaller domain gap and a more automatic workflow than traditional simulation [7]. [8] modeled Lidar effects like secondary returns and ray dropping in the neural rending framework. [9, 10] also leveraged image sequences for additional information. [11, 12] explicitly the temporal dimension to account for the moving objects. [13] rendered both Lidar scans and camera images from the neural field. These approaches realized the playback of a real-world scenario with changed viewpoints and object states, but the pose of the observer and the objects cannot deviate too much from reality, or the rendering quality will deteriorate.
@@ II-D Lidar scan modification
The task in this paper, Lidar scan editing with novel object layout, is related to but different from all the tasks discussed above. Our output is connected to a real scene, similar to the NVS task, but we do not rely on reconstruction, which allows more flexibility in the object manipulation. The task is generative but does not build scenes from scratch, emphasizing the controllability of object layouts.
@@ III Problem Formulation
We address the problem of generating realistic synthetic Lidar data for autonomous driving. Rather than generating data completely from scratch, our approach builds upon real-world scenes, and we aim to manipulate the dynamic objects while preserving the background environment.
The static background of a Lidar scene, denoted as , consists of stationary elements such as streets, trees, buildings, and other fixed background. excludes any foreground dynamic objects of interest, for example, vehicles. The foreground objects in the scene are represented by . A typical driving scene, denoted as , includes both background and foreground objects: . Since it is not feasible to sense the entire from a single viewpoint, we define an observation function , which transforms the scene into a partial-view Lidar point cloud: , which represents the real sensor output for an autonomous vehicle.
@@ IV-A Overview
Our scene editing framework is illustrated in Fig. 2. We first introduce the spherical voxel representation (Sec. IV-B), which is the foundation of our entire pipeline for the efficient modeling of the Lidar projective geometry. Then, we present the three components in our framework. The first involves removing unwanted objects from the Lidar scene (Sec. IV-C). Given an object to be removed, we identify the voxels occupied by the object’s point cloud. We mask these voxels and all voxels occluded by them with an inpainting mask. We remove points in the masked voxels and use a generative point cloud inpainting model to fill the background.
In the second stage, we collect an object library composed of dense full-shape point clouds for every object in all Lidar scans (Sec. IV-D), so that we have a wide range of objects to be inserted back into the Lidar scans with flexible poses. We use the segmentation masks within each Lidar scene to extract partial-view point clouds of vehicles. These partial views are then transformed into full point clouds using a learning-based point cloud completion method.
In the third stage, we insert new objects into the Lidar point cloud (Sec. IV-E). We start by selecting an object from the full-view object library, and specifying a desired pose for its placement. The dense point cloud of this object is then positioned on the ground at the specified pose. The resampling of the inserted objects and the occlusion are then both handled in the spherical voxel representation.
It is worth noting that in our framework, moving an object to a different pose is treated as both removing and re-adding the same object. This approach helps resolve any gaps in the background and allows for significant pose changes.
@@ IV-B Spherical Voxelization
The spherical voxel representation is core to our entire pipeline for efficient occlusion handling and density control, as illustrated in Fig. 3. Spherical voxelization discretizes the space around the origin based on three parameters : radius , azimuth , and elevation . This approach closely mirrors how a LiDAR scan operates, where rays are emitted from the sensor (origin) at specific angles, allowing point resampling with a pattern consistent with real LiDAR sensors. Voxels sharing the same azimuth and elevation values resemble a single LiDAR ray, embedding the occlusion relationship into the data structure (see Fig. 3-right). While the range-view representation could also model the occlusion, the 2D projection loses the 3D structure. Preserving the 3D structure allows effective feature learning and convenient point cloud manipulation in our pipeline.
@@ IV-C Object removal
The object removal has two steps: identifying the area affected by the removal of an object and generating background points in this area that are consistent with the surrounding environment. We call these two steps de-occlusion masking and background inpainting, respectively.
@@ IV-C1 De-occlusion masking
Given the spherical voxelization of the input, a de-occlusion mask refers to a binary mask over the voxels, where the voxels occupied or occluded by the object to be removed are labeled positive. The occupancy of the masked voxels needs to be predicted because of the scene change (object removed) and the visibility change (occlusion removed). We first mask the voxels occupied by the object point cloud. Then, all voxels sharing the same azimuth and elevation indices with larger radius indices than the occupied voxels are also masked (see Fig. 3-right).
@@ IV-C2 Background inpainting
The background inpainting can be viewed as a conditional generation problem. We recast the UltraLidar framework [16] for this task. UltraLidar can be viewed as a Lidar version of MaskGIT [17], which is a generative model for images with transformer-based multi-step autoregression in a discrete latent space. The training has two stages, as depicted in Fig. 4. First, we train a VQ-VAE [23] to encode raw point clouds into a latent feature map with discrete latent codes. We use a bird’s-eye-view (BEV) 2D latent map to represent the encoded point cloud. Notice that our BEV uses the azimuth-radius coordinate from the spherical voxelization, meaning that the voxels along the elevation axis are compressed to a single point in the latent feature map. It is slightly different from the physical top-down projection and also different from the Cartesian voxelization used in UltraLidar. We use the original point cloud scans to learn this encoder.
The second step is to learn a generative model in the latent space for background inpainting. Similar to MaskGIT [17], a bi-directional transformer is trained to map an incomplete BEV latent feature map to a complete latent feature map, which can be decoded to a point cloud. The feature prediction is cast as a classification problem because the feature space is discrete. The prediction is iterative. We use a mask to keep track of which tokens in the BEV feature map need to be predicted. In every iteration, the transformer predicts the masked tokens from the known ones, but only a subset of the predicted tokens with the top classification confidence is preserved, and they will be treated as known tokens in the next iteration. The other tokens remain masked and need to be predicted in the following iterations until the feature map is complete. The BEV mask is initialized from the voxel mask that is used to remove points from the input point cloud. In training, we apply the voxel mask on random object-free voxels, so that the transformer is trained to predict only the background for the masked tokens. At inference, we use the de-occlusion voxel mask introduced in Sec. IV-C1.
While we follow the UltraLidar framework [16], the object removal in UltraLidar is realized by replacing the token at the object position with a background discrete feature, which needs to be manually specified and does not explicitly inpaint the areas occluded by the removed objects. Our object-removal module resolves both problems.
@@ IV-D Object library construction
We create a library of objects with full 3D shapes to be inserted at arbitrary poses in a Lidar scene. The objects are collected from all Lidar scans in our training data. We choose the point cloud as the shape representation of objects for two reasons. First, we can resample points from the shape conveniently in the spherical voxel representation and do not need expensive ray-casting on meshes or volumetric rendering on implicit fields. Second, point cloud completion techniques are well-established to facilitate full-shape generation. We applied a pretrained model of the state-of-the-art point cloud completion network, AnchorFormer [24], in our framework.
@@ IV-E1 Object placement
We assume that a user only specifies the x-y coordinate and the yaw angle of the desired insertion pose, and the insertion algorithm needs to figure out the rest of the degrees of freedom based on the existing background point cloud. We further assume zero pitch and roll angles as are commonly assumed in self-driving datasets [25, 26]. Therefore, we are left with height to be determined.
@@ IV-E2 Resampling and resolving occlusion
These two problems can be efficiently fixed with our spherical voxelization technique. We locate the voxels occupied by the point cloud of the inserted vehicle. Then, we pick the center of each occupied voxel as the resampled points for the shape. Now, the inserted object has a pattern and density that are consistent with the overall Lidar scan. In the end, we resolve the occlusion. For each coordinate with at least one occupied voxel, an occlusion mask is applied to all voxels with radius , where is the smallest radius of the occupied voxels in the 1D ray. All the masked voxels are set to empty.
@@ V Experiments and results
This paper aims at a novel Lidar editing task that is different from the literature. Therefore, it is not straightforward to apply a well-established evaluation protocol and compare with existing work. We will separate the experiments into two major components, i.e., object removal and object insertion, so that we can validate the effectiveness of each step and show that the overall framework provides a practical solution for Lidar scene editing.
We conduct experiments on the nuScenes-LidarSeg [25] dataset, which is composed of 40,000 point clouds sampled from 1,000 driving scenes. The point clouds are collected by a 32-beam Velodyne HDL32E Lidar. We use the official training and validation splits. All results are trained on the Lidar scans in the training set and evaluated using scans in the validation set. We focus on the manipulation of vehicle-type objects, including cars, trucks, and buses. The specific use of the dataset in each experiment is explained in the regarding sections.
@@ V-A Object removal and background inpainting
The goal of this experiment is to test that our approach can inpaint the background points when objects are removed. The tricky thing is that we do not have the ground truth background behind an object. Therefore, we create artificial occlusion masks over the background area in a Lidar scan and ask our model to recover the masked area. In this way, we have the ground truth background to facilitate evaluation.
In order to make the artificial masks similar to the ones caused by the occlusion of real objects, we create the masks from the bounding boxes of objects in nuScenes. This process is illustrated in Fig. 5. We calculate the average size of the 3D bounding boxes of all vehicles in nuScenes. Then, we put the bounding box at a fixed distance, with the heading orthogonal to the viewing direction, so that the box occludes the most azimuth range. We rotate the bounding box around the ego vehicle till the bounding box falls into an object-free sector of the Lidar scan to be masked. We pick the distance as 10 meters to create a substantial occlusion.
We evaluate the performance of our background inpainting using statistical metrics and perceptual metrics. For the statistical metrics, we leverage the Maximum-Mean Discrepancy (MMD) and the Jensen-Shannon Divergence (JSD). Similar to [16], we construct a 2D histogram along the ground plane and use the voxel occupancy instead of the number of points for the bin count. However, our evaluation has two differences from that in [16]. First, our 2D histogram is in the azimuth-radius coordinate. Second, we manually set the count for all bins that are not in the generated area to zero and re-normalize the histogram to sum to one, because we only generate the point cloud for a masked area instead of generating the whole point cloud from scratch. For the perceptual metrics, we use the Frechet Sparse Volume Distance (FSVD) and the Frechet Point-based Volume Distance (FPVD). Following [15], FSVD is evaluated using MinkowskiNet [27], and FPVD is evaluated using SPVCNN [28], both pretrained on nuScenes.
@@
| Metrics | mAP | Car | Truck | Bus |
| Test on nuScenes | 0.551 | 0.704 | 0.386 | 0.564 |
| Test on our data | 0.383 | 0.762 | 0.164 | 0.222 |
| Metrics | mAP | Car | Truck | Bus |
| Without our data | 0.417 | 0.627 | 0.292 | 0.333 |
| Concatenated with our data | 0.432 | 0.631 | 0.285 | 0.381 |
We are not aware of any previous work that fulfills the exact same task. Therefore, we compare our method to a naive inpainting baseline, which is to copy the neighboring object-free point cloud sector and paste and tile it in the masked area. The evaluation results are shown in Tab. I. Our method outperforms the baseline in all metrics, showing the effectiveness of our framework. Additionally, qualitative comparisons are shown in Fig. 6, where it is evident that our model inpaints the background based on the context, resulting in point clouds that appear natural and realistic.

### v2

# LiDAR-EDIT: LiDAR Data Generation by Editing the Object Layouts in Real-World Scenes

@@ Abstract
We present LiDAR-EDIT, a novel paradigm for generating synthetic LiDAR data for autonomous driving. Our framework edits real-world LiDAR scans by introducing new object layouts while preserving the realism of the background environment. Compared to end-to-end frameworks that generate LiDAR point clouds from scratch, LiDAR-EDIT offers users full control over the object layout, including the number, type, and pose of objects, while keeping most of the original real-world background. Our method also provides object labels for the generated data. Compared to novel view synthesis techniques, our framework allows for the creation of counterfactual scenarios with object layouts significantly different from the original real-world scene. LiDAR-EDIT uses spherical voxelization to enforce correct LiDAR projective geometry in the generated point clouds by construction. During object removal and insertion, generative models are employed to fill the unseen background and object parts that were occluded in the original real LiDAR scans. Experimental results demonstrate that our framework produces realistic LiDAR scans with practical value for downstream tasks. Project website with open-sourced code: https://sites.google.com/view/lidar-edit
@@ I Introduction
Data plays a crucial role in the advancement of self-driving technology. High-quality data that accurately captures the complexity and diversity of real-world driving scenes is essential for both the development and testing of autonomous systems. However, real-world data collection is inherently limited due to its high cost and potential safety hazards, especially when capturing long-tail corner cases, which occur infrequently and involve significant risk. Furthermore, real-world data simply represents a fixed recording of past events, making it inflexible for modifying specific factors to analyze their causal impact on self-driving algorithm performance.
These challenges have motivated research focused on generating synthetic data for autonomous driving. Early efforts rely on gaming engines and graphical rendering. This approach generates data from highly configurable virtual environments, but it induces a non-negligible sim-to-real gap, and the scalability is limited. Recent advancements in generative models [1, 2], have enabled end-to-end data generation [3, 4], making large-scale and realistic synthetic data possible. However, this approach comes with a significant drawback: users have limited control over the generated data, such as precisely specifying where a newly generated car will be positioned within the scene. Another notable trend is novel-view synthesis from real-world sequential observations [5], allowing the rendering of a recorded scenario from different viewpoints. However, the variation flexibility of the rendered data is limited by the visible signal.
In this paper, we propose a new paradigm for synthetic data generation: generative 3D scene editing with controllable novel object layouts. We focus on LiDAR scans—a widely used and reliable sensory modality in autonomous vehicles. This paradigm strikes a balance between three key factors: faithfulness, controllability, and scalability. Faithfulness: the generated data are only partial modifications of the real-world data, preserving the characteristics of the real environment and objects in the original data, thereby reducing the synthetic-to-real domain gap. Controllability: users have full control over the object layout in the generated data, including the number, type, and pose of objects, which better facilitates comparative analysis. Scalability: our paradigm does not rely on a physics engine, prebuilt 3D assets, or reconstruction of the whole scene. Instead, it is generative, facilitating efficient and scalable data creation based on vast amounts of existing real-world data.
An illustration of our proposed method for this paradigm, LiDAR-EDIT, is shown in Fig. 1. Our framework first removes unwanted objects from the scene and inpaints missing regions using a generative model. We then insert the point clouds of objects, which are completed from their partial LiDAR scan, at user-specified poses. We use spherical voxelization to ensure that the occlusion and the scanning pattern follow the LiDAR projective geometry. Experimental results demonstrate that our framework enables flexible, controllable, and realistic LiDAR scene editing, offering significant benefits for both the development and testing of autonomous driving algorithms.
@@ II Related Work
We review existing paradigms for LiDAR data generation in the literature.
@@ II-A LiDAR simulation
LiDAR simulation creates a 3D environment with object assets and generates point clouds via ray-surface intersection. [6]. Graphics-based engines [7, 8, 9] face scalability limits from pre-built 3D asset requirements, while asset quality and rendering contribute to domain gaps. [10] used real-world scans to generate 3D assets and applied learning-based postprocessing to reduce this gap, while [11] explored LiDAR phenomena to enhance simulator fidelity.
@@ II-B LiDAR novel-view synthesis
LiDAR novel-view synthesis (NVS) and LiDAR simulation follow the same 3D construction and rendering process, but modern LiDAR NVS integrates both steps into a unified learnable framework called neural rendering. This approach optimizes a learned implicit geometry-appearance model through differentiable rendering for more realistic synthesis, reducing the domain gap and enabling a more automated workflow than traditional simulation [12]. [5] incorporated LiDAR effects like secondary returns and ray dropping, while [13, 14] used image sequences for additional data. [15, 16] accounted for moving objects by modeling the temporal dimension. [17] rendered both LiDAR scans and camera images from the neural field. These methods enable real-world scenario playback with altered viewpoints and object states, though significant deviations in the observer or object poses degrade rendering quality.
@@ II-D LiDAR scan modification
The task in this paper, LiDAR scan editing with novel object layout, is related to but different from all the tasks discussed above. Our output is connected to a real scene, similar to the NVS task, but we do not rely on reconstruction, which allows more flexibility in object manipulation. The task is generative but does not build scenes from scratch, emphasizing the controllability of object layouts.
@@ III Problem Formulation
We address the problem of generating realistic synthetic LiDAR data for autonomous driving. Rather than generating data from scratch, we build upon real-world scenes to manipulate foreground objects while preserving the background environment.
The static background of a LiDAR scene, , consists of stationary elements like streets, trees, and buildings, excluding foreground dynamic objects of interest such as vehicles, represented by . A typical driving scene includes both: . Since sensing the entire from a single viewpoint is infeasible, we define an observation function that transforms the scene into a partial-view point cloud: , representing the real sensor output.
@@ IV-A Overview
Our scene editing framework LiDAR-EDIT is illustrated in Fig. 2. We first introduce the spherical voxel representation (Sec. IV-B), which forms the foundation of our pipeline for efficiently modeling LiDAR projective geometry. Then, we present the three core components of LiDAR-EDIT. The first removes unwanted objects from the LiDAR scene (Sec. IV-C) by identifying voxels occupied by the object and voxels occluded by them. We mask these voxels and use a generative inpainting model to fill the background.
In the second stage, we build an object library of dense full-shape point clouds (Sec. IV-D) for insertion with flexible poses. We extract partial-view point clouds using segmentation masks and transform them into full point clouds via learning-based completion.
In the third stage, we insert objects into the LiDAR point cloud (Sec. IV-E) by selecting an object from our library and specifying a desired pose. The object is positioned on the ground, with resampling and occlusion handled through spherical voxelization.
In our framework, moving an object is implemented as removing and re-adding it, which resolves background gaps and enables significant pose changes.
@@ IV-B Spherical Voxelization
The spherical voxel representation is core to our pipeline for efficient occlusion handling and density control (Fig. 3). Spherical voxelization discretizes space around the origin using three parameters: radius , azimuth , and elevation . This mirrors LiDAR operation, where rays emit from the sensor at specific angles, allowing point resampling with a pattern consistent with real LiDAR sensors. Voxels sharing azimuth and elevation values form a single LiDAR ray, embedding occlusion relationships directly (Fig. 3-right). While range-view representation could model occlusion, it loses 3D structure in 2D projection. Preserving this 3D structure enables effective feature learning and point cloud manipulation.
@@ IV-C Object removal
The object removal stage involves two steps: identifying the area affected by object removal and generating consistent background points in this area. We term these steps de-occlusion masking and background inpainting.
@@ IV-C1 De-occlusion masking
Given spherical voxelization and the object to be removed, the de-occlusion mask highlights voxels whose occupancy needs prediction after removal—specifically, voxels occupied or occluded by the object. We first mask voxels occupied by the object, then mask all voxels sharing the same azimuth and elevation with larger radius indices than the occupied voxels (see Fig. 3-right).
@@ IV-C2 Background inpainting
We approach background inpainting as a conditional generation problem using a method inspired by the UltraLidar framework [19], which is a LiDAR adaptation of MaskGIT[20] that employs transformer-based multi-step autoregression in discrete latent space. The training has two stages, as depicted in Fig. 4. First, we train a VQ-VAE [26] to encode raw point clouds into a discrete latent feature map. We represent this using a bird’s-eye-view (BEV) 2D map with azimuth-radius coordinates, compressing elevation-angle-axis voxels to single latent points. This differs from physical top-down projection and UltraLidar’s Cartesian voxelization. We use the original point cloud scans to train this encoder.
The second step learns a generative model in latent space for background inpainting. Following MaskGIT [20], we train a bi-directional transformer to map incomplete BEV feature maps to complete ones, which decode to point clouds. Feature prediction is treated as a classification problem of the discrete latent codes with iterative processing. A mask tracks which tokens need prediction, with each iteration preserving only sampled high-confidence predictions as known tokens for subsequent iterations. The BEV mask initializes from the voxel mask used to remove points. During training, we apply the mask to random object-free voxels, training the transformer to predict only background for masked tokens. At inference, we use the de-occlusion mask from Sec. IV-C1.
LiDAR-EDIT inherits UltraLidar’s two-stage framework [19], while tailoring the system to generate realistic backgrounds after object removal by locating occluded areas via spherical voxelization and using background-only training. In comparison, UltraLidar simply replaces object tokens with manually-specified discrete background features without guaranteeing complete background inpainting.
@@ IV-D Object library construction
We create a library of objects with full 3D shapes to be inserted at arbitrary poses in a LiDAR scene. The objects are collected from all LiDAR scans in our training data. We choose the point cloud as the shape representation of objects for two reasons. First, we can resample points from the shape conveniently through spherical voxelization without ray-casting on meshes or volumetric rendering on implicit fields. Second, point cloud completion techniques are well-established to facilitate full-shape generation. We apply a pretrained model of the state-of-the-art point cloud completion network, AnchorFormer [27], in our framework.
@@ IV-E1 Object placement
We assume that a user only specifies the x-y coordinate and the yaw angle of the desired insertion pose, and the insertion algorithm needs to figure out the rest of the degrees of freedom based on the existing background point cloud. We further assume zero pitch and roll angles, as commonly assumed in self-driving datasets [28, 29]. Therefore, we are left with height to be determined.
@@ IV-E2 Resampling and resolving occlusion
These two problems can be efficiently fixed with our spherical voxelization technique. We locate the voxels occupied by the point cloud of the inserted vehicle. Then, we pick the center of each occupied voxel as the resampled points for the shape. Now, the inserted object has a pattern and density that are consistent with the overall LiDAR scan. In the end, we resolve all occlusions using spherical projection: for each coordinate with at least one occupied voxel, an occlusion mask is applied to all voxels with radius , where is the smallest radius of the occupied voxels in the 1D ray. All the masked voxels are set to empty.
@@ V Experiments and results
This paper introduces a novel LiDAR editing approach distinct from all existing methods. Therefore, it is not straightforward to use a well-established evaluation protocol or compare LiDAR-EDIT with previous works. We divide the experiments into two major components, object removal and object insertion, showing our framework generates realistic LiDAR scenes.
We conduct experiments on the nuScenes-LidarSeg [28] dataset, consisting of 40,000 point clouds from 1,000 driving scenes, captured by a 32-beam Velodyne HDL32E LiDAR. Using the official training and validation splits, all models are trained on the training set of size 28,130, and evaluated on the validation set of size 6,019. The dataset’s specific use in each experiment is detailed in the relevant sections. In all experiments, we use spherical voxelization of resolution within the range of along the radius, azimuth and elevation axes respectively.
@@ V-A Object removal and background inpainting
We design two experiments to evaluate the quality of our inpainted background and to verify that our pipeline can generate object-free point clouds after object removal and background inpainting, respectively.
The first experiment evaluates LiDAR-EDIT’s ability to inpaint realistic background after vehicle removal. Since ground truth background behind real objects is not available, we create artificial de-occlusion masks over background areas and task our model with reconstructing these masked points, giving us ground-truth for evaluation.
To create masks resembling real object occlusions, we use bounding boxes from nuScenes objects (Fig. 5). We calculate the average size of all ground truth vehicle bounding boxes, then place a box of the average size at a fixed 10-meter distance with heading orthogonal to viewing direction to maximize azimuth occlusion. We rotate this box until it falls within an object-free sector of the LiDAR scan.
We evaluate our generated background against ground truth using both statistical and perceptual metrics. Statistical metrics quantifythe global structural differences between the two sets of background poins, while perceptual metrics assess differences in the latent feature space, capturing both global and local structures. For statistical metrics, we use Maximum-Mean Discrepancy (MMD) and Jensen-Shannon Divergence (JSD). Following [19], we construct a 2D histogram along the ground plane, using voxel occupancy instead of point count for bin values. Our evaluation differs in two ways: (1) our 2D histogram is in azimuth-radius coordinates, and (2) we zero-out bins outside the generated area and re-normalize the histogram, as we generate only a masked region rather than an entire point cloud. For perceptual metrics, we use Frechet Sparse Volume Distance (FSVD) and Frechet Point-based Volume Distance (FPVD). Following [18], FSVD is evaluated using MinkowskiNet [30], and FPVD is evaluated using SPVCNN [31], both pretrained on nuScenes.
@@
| Data generation method | Ours | UltraLidar [19] |
| Number of objects detected per frame | 2.6 | 15.5 |
As no prior work addresses this exact task, we compare our method to a naive inpainting baseline, which copies and tiles a neighboring object-free point cloud sector into the masked area. Using the nuScenes validation set for evaluation, Tab. I shows that our method outperforms the baseline across all metrics, demonstrating its effectiveness. Fig. 6 provides qualitative comparisons, highlighting how our model inpaints the background based on the context, producing more natural and realistic point clouds.
In the second experiment, we use de-occlusion masks from ground-truth object bounding boxes to generate object-free point clouds. We then use an object detection model, VoxelNext [32], to count objects in the point clouds, and the results are in Tab. II. Ideally, the count should be zero; however, a small nonzero value is observed due to false positives, which is expected. For reference, UltraLidar’s unconditionally generated point clouds, which are expected to contain objects, yield an average of 15.5 detected objects.
| Training data | Testing data | mAP | car | non-car |
| nuScenes | Ours | 0.340 | 0.785 | 0.290 |
| nuScenes | nuScenes | 0.361 | 0.672 | 0.326 |
| Ours + nuScenes | nuScenes | 0.368 | 0.672 | 0.334 |